OpenStudy (anonymous):

Suppose you draw a card from a deck of 52 cards. What is the probability that the card you draw is NOT an ace? There are 4 aces in the deck.

OpenStudy (anonymous):

An Ace: There are 4 aces out of 52 cards, so the probability is P(A) = 4/52 = 1/13 so.... just reverse it... take out the four aces how many cards other then the aces do u have, over all the cards u have.. simplify that after then if u can

OpenStudy (cunningham2000):

48/52 = 92.3% probability or 12/13

OpenStudy (anonymous):

yay.. :) theres the answer!!!

OpenStudy (anonymous):

thanks guys :)